# FORWARD: Dataset of a forwarder operating in rough terrain

### v1

@@ 6.2 Vibration data
For a subset of the data, we have vibration data recorded from the operator seat. The vibration data is recorded at 6 000 Hz and the vibration dose value (VDV), whole-body vibration exposure A(8), and daily equivalent static compression dose Sed is computed for a number of test scenarios and regular forwarding work (Table 4). Vibration dose value (VDV) and whole body vibration exposure A(8) is computed from the raw data using the ISO 2631-1 ([15]) standard, while a daily equivalent static compression dose (Sed) value is computed according to ISO 2631-5 ([16]). For a stitched scenario containing 11 hours and 20 minutes of regular forwarding work, the VDV is 20.0 m/s1.75 and A(8) is 0.66 m/s2. We also synchronize the data from the vibration sensor with the machine-data by downsampling it to 5 Hz, using metrics such as maximum, minimum, average, and standard deviation. The downsampled vibration data lack sufficient resolution to be used for computation of human vibration exposure measures, but it is useful for exploring similarities between the vibration data and the machine data.
@@
Using the maximum value for each data point as downsampling statistic, we plot the downsampled vibration data together with the machine accelerometer data for a subset of the test circuit covering the passage of a 90 cm high rock (Fig 19). In the Z-direction, peaks in acceleration are observed when the forwarder drives over obstacles, such as rocks and stumps. In this specific case, the vibration data shows a clear peak in the Z-direction when the forwarder drives down from the 90 cm high rock (Fig 19).
Figure 19: Downsampled machine vibration data (left) and operator seat vibration data (right) during the passage of a 90 cm high rock on the test circuit. Plotted along the machine path with high values in red and low values in blue and as a time series. Both datasets are accelerations in the Z direction.
@@ 6.3 StanForD production data
7 Conclusions
The FORWARD dataset is unique in the forestry sector with its high temporal and spatial resolution, and the combination of machine data, external sensor data, terrain data, and video data. Statistics about e.g., machine utilization and performance from long time-series are traditionally being collected in Sweden, but these datasets are not sufficient in resolution for the needs of modern forestry operations. To study and improve machine performance, energy use, machine design, terrain-machine interactions, and operator comfort in detail, high-resolution data from both terrain and machine is needed. The FORWARD dataset fills this gap and provides a valuable resource for researchers and practitioners in the fields of forestry operations, physics-based simulations, machine learning, and robotics. The dataset is curated and synchronized, making it easy to use for a variety of analyses and algorithm training purposes.
Our philosophy when compiling the dataset has been to include as much data and code as possible, and let the users of the dataset decide what is relevant for their specific use case. We therefore encourage users to take advantage of the provided analysis scripts and extend them with their own code to fit scientific or practical research projects that focus beyond our initial motivation to gather the data.
8 Acknowledgements
@@
This work was supported in part by Mistra Digital Forest (Grant DIA 2017/14 #6) and Horizon Europe Project XSCAVE under Grant 101189836. The authors would like to thank the Swedish forest company SCA and machine manufacturer Komatsu Forest for providing the ALS- and machine data used in this study via collaborative field studies. We also thank Christian Höök, SLU, for his generous help with video data collection 2023. Hans Pettersson, Umeå University, is acknowledged for his help with the vibration data collection and processing.

### v2

@@ 4.4 Vibration data
For a subset of the data, we have vibration data recorded from the operator seat. The vibration data is recorded at 6 000 Hz and the vibration dose value (VDV), whole-body vibration exposure A(8), and daily equivalent static compression dose Sed is computed for a number of test scenarios and regular forwarding work (Table 4). Vibration dose value (VDV) and whole body vibration exposure A(8) is computed from the raw data using the ISO 2631-1 standard [undefn], while a daily equivalent static compression dose (Sed) value is computed according to ISO 2631-5 [undefo]. For a stitched scenario containing 11 hours and 20 minutes of regular forwarding work, the VDV is 20.0 m/s1.75 and A(8) is 0.66 m/s2. We also synchronize the data from the vibration sensor with the machine-data by downsampling it to 5 Hz, using metrics such as maximum, minimum, average, and standard deviation. The downsampled vibration data lack sufficient resolution to be used for computation of human vibration exposure measures, but it is useful for exploring similarities between the vibration data and the machine data.
@@
Using the maximum value for each data point as downsampling statistic, we plot the downsampled vibration data together with the machine accelerometer data for a subset of the test circuit covering the passage of a 90 cm high rock (Fig 19). In the vertical direction, peaks in acceleration are observed when the forwarder drives over obstacles, such as rocks and stumps. In this specific case, the vibration data shows a clear peak in the vertical direction when the forwarder drives down from the 90 cm high rock (Fig 19).
Figure 19: Downsampled machine vibration data (left) and operator seat vibration data (right) during the passage of a 90 cm high rock on the test circuit. Plotted along the machine path with high values in red and low values in blue and as a time series. Both datasets are accelerations in the vertical direction.
@@ 4.5 StanForD production data
Limitations
Some notes can be made about expected accuracy for different parts of the dataset:
  • After further analysis of the driving speed it has became clear that there are more settings in the machine than Throttle position, Inch factor, and Gear that have effect on the target driving speed of the forwarder. There is a variable called transmission mode that controls the diesel engine rpm in different situations and thus the target driving speed, however this variable is not included in the dataset due to confidentiality reasons.
  • GNSS-variables from the machine data, such as positions, heading, and speed are captured with a system developed by Komatsu Forest that use two RTK-GNSS recievers on the cabin roof. When internet connection is available (most of the time) the system is connected to the Swedish SWEPOS network of base stations, enabling centimeter-level accuracy for positions. The dual reciever setup enable computation of heading. The accuracy variable value ’RTK_Fix’ in the dataset indicate optimal accuracy for the position and heading respectively.
  • The video files are cut or muted to remove conversations held between tests.
  • For pointcloud data, generated digital terrain models, and plotted machine paths, the coordinate reference system SWEREF99 - TM (EPSG:3006) is used. The experimental sites are situated relatively close to Sweden’s central meridiane, mitigating east-west related distortions.
  • Drone orthomosaics from the Märrviken site have been manually corrected in horizontal directions by aligning visible objects with the terrain models produced from helicopter-borne LiDAR. This was necessary since the drone used 2023 did not have any positional correction available and no manual control points were arranged in the field. The level of accuracy is sufficient for using the orthomosaic for visualization purposes together with the LiDAR data, for other uses it may be insufficient.
Ethics statement
The authors confirm that they have read and followed the ethical requirements for publication in Data in Brief. The current work does not involve human participants, animal experiments, or any data collected from social media platforms.
Credit author statement
Mikael Lundbäck: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Erik Wallin: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization. Carola Häggström: Conceptualization, Methodology, Validation, Investigation, Resources, Data Curation, Writing - Review & Editing, Supervision. Mattias Nyström: Conceptualization, Methodology, Validation, Investigation, Resources, Data Curation, Writing - Review & Editing. Andreas Grönlund: Methodology, Validation, Resources, Writing - Review & Editing, Project administration. Mats Richardson: Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Review & Editing. Petrus Jönsson: Conceptualization, Methodology, Investigation, Resources, Writing - Review & Editing, Project administration, Funding acquisition. William Arnvik: Formal analysis, Investigation, Data Curation, Writing - Review & Editing, Visualization. Lucas Hedström: Software, Validation, Formal analysis, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization. Arvid Fälldin: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Martin Servin: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Resources, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision, Project administration, Funding acquisition.
Acknowledgements
@@
This work was supported in part by Mistra Digital Forest (Grant DIA 2017/14 #6) and Horizon Europe Project XSCAVE under Grant 101189836. The authors would like to thank the forest company SCA for providing the aerial LiDAR data and machine- and personnel resources used in this study via collaborative field studies. We also thank Christian Höök, SLU, for his generous help with video data collection 2023. Hans Pettersson, Umeå University, is acknowledged for his help with the vibration data collection and processing.
Declaration of competing interest
Mattias Nyström and Andreas Grönlund are employed by Komatsu Forest AB, Umeå, Sweden. The authors declare that the research was conducted in the absence of any commercial, personal or financial relationships that could be construed as a potential conflict of interest.